OpenStudy (anonymous):

Find the solution to the system of the following equations. You may use either the graphing or substitution method. y=x+3 y=2x+4 -1,2 2,-1 0,3 0,4

OpenStudy (anonymous):

It would be (-1,2) I pluged it into the equation 2=2(-1)+4 2=-2+4 2=2

OpenStudy (anonymous):

And y=x+3 2=-1+3 2=2

OpenStudy (kirbykirby):

If you want an algebraic way, by substitution, put in what y equals into the second equation: \[ x+3=2x+4\\ x-2x=4-3\\ -x=1\\ x=-1\] Then plug in this x value into say the 1st equation: \[ y=(-1)+3\\ y=2\]